1 mole of non ideal gas at 25*C is subjected to expand reversibly ten times of the initial volume. The change in entropy of expansion is,
1
Expert's answer
2020-06-08T15:46:02-0400

We have the state relation -

"TdS=dU+PdV"

R=8.314Jmol−1

"\u200b\\frac{V2}{V1}" =10

Since expansion is isothermal, there will be no change in temperature and consequently no change in internal energy ("dU = 0" ).

It reduces to -

"ds=\frac{PdV}{T}"

From ideal gas relation, "Pv= nRT"

"dS=\\frac{nRdV}{V}\n\u200b"

"S 2\u2212S1 =nRln( \\frac{V1}{V2}\n\u200b\n\u200b\t\n \n\u200b\t\n )"

n = 1 (1 mole of gas), R=8.314Jmol-1K,

S2−S1=8.314×ln(10)

⟹S2−S1=19.15K-1mole-1